# Implicit Actor Critic Coupling via a Supervised Learning Framework for RLVR

### v1

@@ C.4 Ablation Study
To conduct a comparative analysis of how different advantage estimators affect PACS performance, we benchmark our default RLOO method against two alternatives: GRPO (Shao et al., 2024) and Dr. GRPO (Liu et al., 2025b). While RLOO utilizes a leave-one-out mechanism, the advantage functions for GRPO and Dr. GRPO are defined as follows. Dr. GRPO introduces simple yet significant modifications to address the biases in GRPO by removing the std normalization terms.
@@
Model $\bm{k=1}$ $\bm{2}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$
PACS 78.97 84.39 88.18 90.68 92.33 93.60
    - GRPO 74.71 80.19 84.31 87.42 89.64 91.20
    - Dr. GRPO 79.68 84.44 87.90 90.52 92.30 93.40
Table 11: Results of different advantage estimators of Qwen2.5-7B on MATH 500. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
@@
Model $\bm{k=1}$ $\bm{2}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$
PACS 57.83 65.57 71.94 78.17 83.83 88.22 91.95 95.51 98.51
    - GRPO 48.66 57.43 63.96 69.68 75.71 81.63 86.94 90.95 94.06
    - Dr. GRPO 58.29 66.54 72.83 78.27 83.28 87.45 90.12 91.65 93.13
Table 12: Results of different advantage estimators of Qwen2.5-7B on AMC23. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
On the MATH 500 and AMC23 benchmarks (see Tables 11 and  12), PACS (employing RLOO) demonstrates highly comparable performance with the Dr. GRPO variant. This result aligns with expectations, as RLOO and Dr. GRPO exhibit only subtle algorithmic differences, both enhancing the foundational GRPO through intra-batch relative comparisons. On these moderately challenging tasks, such minor algorithmic distinctions do not translate into substantial performance disparities. Specifically, Dr. GRPO shows marginal superiority at lower $k$ (kk=1, 2), while PACS exhibits superior performance at higher $k$ values; however, the overall performance of both methods remains remarkably similar.
@@
Model $\bm{k=1}$ $\bm{2}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$
PACS 15.80 19.93 24.39 29.75 36.29 43.90 51.21 56.54 59.78
    - GRPO 12.75 16.24 19.00 21.59 24.88 28.91 33.22 37.14 39.95
    - Dr. GRPO 15.59 19.67 23.99 29.04 34.78 40.82 46.76 53.08 59.68
Table 13: Results of different advantage estimators of Qwen2.5-7B on AIME 2024. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
@@
Model $\bm{k=1}$ $\bm{2}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$ $\bm{128}$ $\bm{256}$
PACS 13.33 18.19 23.73 29.65 36.00 42.88 49.45 54.48 58.22
    - GRPO 8.39 13.01 17.69 21.78 25.77 29.86 34.34 39.26 43.73
    - Dr. GRPO 13.93 17.15 20.23 24.05 28.68 34.55 41.07 46.12 49.56
Table 14: Results of different advantage estimators of Qwen2.5-7B on AIME 2025. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
The superiority of RLOO is most pronounced on the AIME 24 and AIME 25 (Tables 13 and 14). Here, the nuanced algorithmic difference between RLOO and Dr. GRPO is amplified into a clear performance divergence, with our PACS algorithm demonstrating unequivocal superiority across all values of $k$. We posit that this performance gap is rooted in the distinctive ”leave-one-out” credit assignment of RLOO. High-difficulty tasks with sparse rewards and high solution variance demand a robust signal for policy updates. RLOO’s mechanism—contrasting each sample against the cohort average—provides exactly this stability. The practical consequence is substantial: on the AIME 2025 benchmark, for instance, PACS (58.22%) establishes a lead of almost 9 percentage points over Dr. GRPO (49.56%) at pass@256.

### v2

@@ D.3 Ablation of Different Advantage Estimators
To conduct a comparative analysis of how different advantage estimators affect PACS performance, we benchmark our default RLOO method against two alternatives: GRPO (Shao et al., 2024) and Dr. GRPO (Liu et al., 2025c). While RLOO utilizes a leave-one-out mechanism, the advantage functions for GRPO and Dr. GRPO are defined as follows. Dr. GRPO introduces simple yet significant modifications to address the biases in GRPO by removing the std normalization terms.
@@
Model $\bm{k=1}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$
PACS 94.80 97.87 98.47 98.93 99.30 99.49
   - GRPO 94.58 97.78 98.49 99.00 99.27 99.38
   - Dr. GRPO 94.82 97.86 98.48 98.93 99.27 99.47
Table 9: Results of different advantage estimators of Qwen3-4B on MATH 500. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
@@
Model $\bm{k=1}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$
PACS 90.45 96.01 96.96 97.40 97.50 97.50
   - GRPO 89.55 93.72 95.60 96.79 97.59 98.11
   - Dr. GRPO 90.04 95.50 96.49 97.15 97.47 97.50
Table 10: Results of different advantage estimators of Qwen3-4B on AMC23. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
@@
Model $\bm{k=1}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$
PACS 55.10 72.91 77.93 80.74 82.25 83.13
   - GRPO 56.02 67.74 77.01 81.32 83.02 83.33
   - Dr. GRPO 55.78 71.54 76.28 81.27 82.89 83.32
Table 11: Results of different advantage estimators of Qwen3-4B on AIME2024. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
@@
Model $\bm{k=1}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$
PACS 45.63 61.00 66.18 70.14 72.91 74.80
   - GRPO 44.01 59.66 65.92 70.94 75.04 78.96
   - Dr. GRPO 44.95 60.17 65.90 70.85 75.06 79.03
Table 12: Results of different advantage estimators of Qwen3-4B on AIME2025. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
BeyondAIME ($\text{pass@}k$)
Model $\bm{k=1}$ $\bm{4}$ $\bm{8}$ $\bm{16}$ $\bm{32}$ $\bm{64}$
PACS 27.16 41.66 47.96 53.41 57.91 62.10
   - GRPO 26.41 39.95 45.73 50.54 54.03 56.30
   - Dr. GRPO 26.67 40.70 46.87 52.33 56.73 60.36
Table 13: Results of different advantage estimators of Qwen3-4B on BeyondAIME. Bold numbers indicate the best performance. Underlined numbers indicate the second best.
Our findings indicate that PACS with RLOO exhibits superior performance, particularly on challenging benchamarks. As shown in table 13, on the hardest dataset BeyondAIME, PACS achieves a $\text{pass@}1$ of 27.16% and $\text{pass@}64$ of 62.10%, consistently outperforming Dr. GRPO (26.67% / 60.36%) and GRPO (26.41% / 56.30%) across all $k$ values.
While Dr. GRPO demonstrates competitive performance on standard benchmarks like MATH 500 and at higher sample counts on AIME 2025 , PACS maintains a critical edge in the low-sample regime ($\text{pass@}1$) across most tasks (e.g., 90.45% vs. 90.04% on AMC 23 ). We posit that this performance advantage stems from the algorithmic property of RLOO: by contrasting a single sample with the leave-one-out average rather than the group mean including itself, RLOO minimizes the bias in advantage estimation, yielding more precise signals for credit assignment in complex reasoning paths.
Appendix E Diversity Score Definition
To rigorously quantify the semantic diversity of the generated reasoning paths, we utilize an embedding-based metric. Let $\mathcal{Q}=\{q_{1},q_{2},\dots,q_{N}\}$ denote the evaluation dataset consisting of $N$ queries. For each query $q_{i}$, the model generates a set of $K$ candidate outputs, denoted as $\mathcal{O}_{i}=\{o_{i,1},o_{i,2},\dots,o_{i,K}\}$.
We employ a pre-trained embedding model $E(\cdot)$ (specifically Qwen3-0.6B-embedding (Zhang et al., 2025) in our experiments) to map each textual output $o_{i,j}$ into a $d$-dimensional dense vector representation:
$\mathbf{v}_{i,j}=E(o_{i,j})\in\mathbb{R}^{d}$ (12)
For a specific query $q_{i}$, the semantic similarity between two output vectors $\mathbf{v}_{i,m}$ and $\mathbf{v}_{i,n}$ is measured using cosine similarity:
$\text{sim}(\mathbf{v}_{i,m},\mathbf{v}_{i,n})=\frac{\mathbf{v}_{i,m}\cdot\mathbf{v}_{i,n}}{\|\mathbf{v}_{i,m}\|\|\mathbf{v}_{i,n}\|}$ (13)
The diversity score for query $q_{i}$, denoted as $Div(q_{i})$, is defined as the average pairwise cosine distance among the generated responses. To ensure an unbiased estimate, we compute the average over all unique pairs $(m,n)$ where $1\leq m<n\leq K$. Since cosine distance is defined as $1-\text{similarity}$, the formulation is as follows:
$Div(q_{i})=1-\underbrace{\frac{1}{\binom{K}{2}}\sum_{1\leq m<n\leq K}\text{sim}(\mathbf{v}_{i,m},\mathbf{v}_{i,n})}_{\text{Average Pairwise Similarity}}$ (14)
Finally, the global diversity metric $S_{div}$ for the model is calculated by aggregating the diversity scores across all queries in the dataset:
$S_{div}=\frac{1}{N}\sum_{i=1}^{N}Div(q_{i})$ (15)
A higher $S_{div}$ value indicates greater semantic variation among the sampled responses, reflecting a broader exploration of the solution space.